Capacitor Circuits I Each of the circuits below cesists of baery and several capaciors Several of the voltapes, charges and capacitances are indicated Use conservation of charpe and the detinition of capaciance to determine as many of the other voltages, charges, and capacitances as possible B 24 6 HC 12V 12 V 12 uC 30 V 18 uC CI E 14 C esa Conmanity Colkgs, Mex AZ 85202 Page 3.6 and Trevor F. Meder
Answer #1

12valt 12 v AT luF 3 voulh ovalt 80 chatq will bel Sqmu om beth on the capa i tout E VO than thi capa ei torn cup of th capqo chaige will b same on both that, change om capaciło c 4 C. Vel tegu acuto3s the catacitourn A i than charga omb) Same om both both wiill bt same 3+6 Noltaga ga 12u VPA = 6 Volt?e =(11+18.+ 6).uc VE361S Vault that ip., voltaga of the batlunycapaeitance CF = 年--丝 24 2.0 RON -> 16 valt 124 wiill bu Samu om both 1C -thon Ceuta cl ton cu of D igl 6 4- B cmd C a in Parallel 80 vedtagu acut oys both will bue same +ham VB Va= (26-2) = 18 Volt. than charge on A = 1x18 4AQlF 30 3ον 2p= 3x13 change on B = 36-13 thon capa citam cwofc 12 1 uF4 chaige on E 1 ch